### v1

@@ 3.2.1 Formula and number of hyperparameters
TGRF leverages the characteristics of the normal distribution [47]. The normal distribution offers the advantage of being able to transform into various shapes with only two hyperparameters. This flexibility enhances the model’s adaptability, allowing adjustments to fit diverse prior knowledge and apply them to reward functions. Furthermore, it alleviates the burden on researchers in tuning hyperparameters, aiding in the swift identification of appropriate hyperparameters within a short time frame.
@@
Figure 4: TGRF. The X-axis denotes the X-position in meters, the Y-axis represents the Y-position in meters, and the Z-axis indicates TGRF value when $w_{TGRF}=1$,$x_{TGRF}=dist(X,Y)$,$\mu_{TGRF}=0$,$\sigma_{TGRF}=2$.
@@
TGRF leverages characteristics of a normal distribution. In Equation (2), TGRF involves three hyperparameters: $w_{TGRF}$, representing the weight of TGRF; $\mu_{TGRF}$, denoting the mean; and $\rho_{TGRF}$, indicating the variance.
It incorporates a single variable $x_{TGRF}$. $C_{norm}$ ensures that TGRF attains a maximum value of 1 irrespective of $\rho_{TGRF}$. This allows the scaling of TGRF solely by $w_{TGRF}$, enabling researchers to intuitively control its maximum value, preventing the freezing robot problem [43], and balancing it with other reward values to attain the desired algorithmic performance ultimately.
$\rho_{TGRF}$ determines the transformability of TGRF. As $\lim_{\sigma_{TGRF}\to\infty}\mathcal{N}(x_{TGRF};\mu_{TGRF},\sigma_{TGRF})$, it takes on a constant form insensitive to changes in $x_{TGRF}$, while as $\lim_{\sigma_{TGRF}\to 0}\mathcal{N}(x_{TGRF};\mu_{TGRF},\sigma_{TGRF})$, it resembles an impulse function. This versatility enables the creation of diverse forms such as constant, linear, non-linear, Gaussian, etc., adaptable to specific objectives.
$\mu_{TGRF}$ determines the position of TGRF’s maximum value. Thus, it allows adjustments of the maximum value position according to the objectives requiring specific values to be reached. The variable $x_{TGRF}$ signifies the variable in N($x_{TGRF}$; $\mu_{TGRF}$,$\rho_{TGRF}$). As a result, it ultimately represents a shape similar to Fig. 4.
@@ 3.2.2 Transformability
Figure 5: Transformability of TGRF. The X-axis denotes the X-position in meters, the Y-axis represents the Y-position in meters, and the Z-axis indicates TGRF value. In (a), $w_{TGRF}=1$,$x_{TGRF}=dist(X,Y)$,$\mu_{TGRF}=0$,$\sigma_{TGRF}=5$. In (b), $w_{TGRF}=1$,$x_{TGRF}=dist(X,Y)$,$\mu_{TGRF}=0$,$\sigma_{TGRF}=5000$.
The reward functions in Section II-B had different formulas and limited flexibility. These constraints lead to complexity in the reward functions and difficulty in determining the optimal performance. Therefore, researchers must redesign and tune them to achieve better performances. However, the TGRF offers versatility in generating various shapes.
Fig. 5 illustrates the creation of different shapes using the same TGRF by simply adjusting $\rho_{TGRF}$. Fig. 5a results in a TGRF that generates a continuous Gaussian distribution, making it suitable for moving objects or humans by applying varying penalties. Moreover, Fig. 5b forms a discrete column like shape. This configuration is suitable for stationary reward functions or objects. This approach allows various shapes via adjusting only a hyperparameter ($\rho_{TGRF}$) without requiring an additional function. This significantly reduces the researchers’ time and effort while enabling finetuning to match the specific characteristics of objects.
@@ 3.2.3 Reward function
Reward $r(s_{t},a_{t})$ is categorized into five types. First, $r_{goal}(s_{t})$ = +10 represents the reward when the robot successfully reaches its destination. Second, $r_{col}(s_{t})$ = -10 serves as a penalty incurred upon colliding with another individual. Third, $r_{disc}(s_{t})$ represents a penalty for entering a danger zone. Fourth, $r_{pot}(s_{t})$ corresponds to the reward/penalty contingent on the change in distance to destination $S_{goal}$. Finally, $r_{pred}(s_{t})$ denotes the penalty invoked when entering a prediction.
Figure 6: TGRF applied to reward functions. The X-axis denotes the X-position in meters, the Y-axis represents the Y-position in meters, and the Z-axis indicates negative reward value. The central cylinder represents $r_{col}$, and the surrounding distribution represents $r_{disc}(s_{t})$. Beyond $d_{disc}$, $r_{disc}(s_{t})$ becomes 0 ($s_{t}\notin S_{\text{danger zone}}$). In (a), $w_{TGRF}=8$,$x_{TGRF}=dist(X,Y)$,$\mu_{TGRF}=0$,$\sigma_{TGRF}=3$. In (b), $w_{TGRF}=10$,$x_{TGRF}=dist(X,Y)$, $\mu_{TGRF}=0$, $\sigma_{TGRF}=1000$.
Penalty $r_{disc}(s_{t})$ is imposed when the robot enters the danger zone ($d_{min}$ is within $d_{disc}$) determined by the nearest human distance, denoted as $d_{min}$. The formula is as follows:
@@
In (3) and Fig. 6a, $r_{disc}(s_{t})$ is designed to prevent collisions with humans, following a gaussian distribution with $\mu_{disc}$=0, whereby smaller values of $d_{min}$ (closer proximity to humans) result in a higher penalty. This design aligns with prior knowledge and exhibiting a more plausible and realistic shape.
$r_{disc}(s_{t})$ can be adjusted using only two hyperparameters. By tuning $\rho_{disc}$, researchers can regulate the breadth of the gaussian penalty concerning the distance between humans and the robot. This enables the robot to react more sensitively or less sensitively to the distance from humans. $w_{disc}$ directly scales the reward function, establishing a linear correlation with $r_{disc}(s_{t})$, thereby enabling the adjustment of the overall reward balance to prioritize driving tasks.
Potential reward $r_{pot}(s_{t})$ represents the reward associated with the potential field, and is defined as follows:
@@
$r_{pot}(s_{t})$ plays a crucial role in guiding a robot toward its destination and mitigating the freezing robot problem [21]. However, high values of $r_{pot}(s_{t})$ can lead to increased collision rates with humans. Therefore, in equation (4), by setting $x_{pot}=\mu_{pot}$, we aimed to maintain a constant TGRF regardless of $\mu_{pot}$ and $\rho_{pot}$. In addition, by applying $w_{pot}$=1.5, we ensured equal weighting for all $\Delta d$, aiming for uniformity across them.
@@
$r_{pred}(s_{t})$ is used only in some models employing trajectory prediction. $r_{pred}(s_{t})$ denotes the penalty value when the robot is positioned in the trajectory of the $i-th$ person. $\mathbbm{1}_{i}^{t+k}$ indicates whether the robot exists at the predicted position of the $i-th$ person at time t+k or not. Thus, $r_{pred}(s_{t})$ takes the smallest penalty among all individuals’ trajectory penalties that the robot takes. In our experiments, to demonstrate the performance enhancement even with different reward functions, we adopted $r_{pred}(s_{t})$ used in the GST + HH Attn from [13]. The final definition of our reward function is as follows:
@@
In summary, the TGRF offers the distinct advantage of intuitively and efficiently modifying the reward function with fewer hyperparameters. This enables the robot to make rational decisions and reduces the time consuming task of fine-tuning.
@@ 4.1 Experimental environment
Figure 7: Comparing a scenario with TGRF and a scenario without TGRF. Yellow circles represent robots, blue circles represent humans within the sensor range, red circles represent humans outside the sensor range, and orange circles in front of the blue circles indicate trajectories predicted by the GST.
We employed a 2D environmental simulator used in previous studies [13]. This simulator features a 12 × 12m space, with a 360° field of view and 5m sensor range for the light detection and ranging (LiDAR). We used a fixed number of humans (20) to represent a crowded setting.
@@
Both humans and robot operated using holonomic kinematics to determine their velocities $(a_{t}=[v_{x},v_{y}]m/s)$ along the x- and y-axes. Holonomic kinematics refers to a state where degrees of freedom can move independently without any constraints. This implies that robots and machines can move without any limitations on their position or orientation. As a result, the action space of a robot is continuous, allowing both robots and humans to immediately achieve their desired speed within a time frame of $\Delta t$, assuming they operate within the maximum speed limit. Therefore, the positions of the humans and robots were continuously updated according to (7).
The robot has attributions such as a size of $\rho=0.3m$ and maximum speed of $v_{max}$ = $1.0m/s$. Humans also have characteristics such as a size ranging from 0.3 to 0.5m and maximum speed varying between 0.5 and 1.5m/s. In addition, the locations and destinations of the robot were randomized, with the destinations not set excessively close. If the robot collides with an individual, reaches its destination, or exceeds the maximum time T, the episode is terminated, leading to the beginning of the subsequent episode.
Random seeding was applied during training, resulting in varying outcomes for each training episode. To handle the varying outcomes, multiple training runs were conducted with a total time step of 2×$10^{7}$ for the DS-RNN and 1×$10^{7}$ for the other algorithms. The learning rate was set to 4×$10^{-5}$ for all policies. Subsequently, we acquired test data from 500 test episodes. The evaluation metrics applied to the test data included the success rate (SR), average navigation time (NT) in seconds, path length (PL) in meters for successful episodes, and intrusion time ratio (ITR).
@@ 4.2.1 Results in different navigation methods and environment
We compared the performance of rewards (from section III-B-3) when the TGRF was applied to the performance without its application (reward function from [13]). The experiment involved the application of five different navigation methods for robots, as mentioned in Section III-A-1. Table I shows the performance when individuals adhere to ORCA, while Table II outlines the performance when adhering to SF. The hyperparameters are set to $w_{disc}=0.25$, $x_{disc}=d_{min}$, $\mu_{disc}=0$,$\rho_{disc}=0.2$, $d_{disc}=0.5$, $w_{pot}=1.5$,$x_{pot}=\mu_{pot}$, $\mu_{pot}=0$,$\rho_{pot}=1000$. Through Table 1, 2, and Fig. 7, we were able to identify the three impacts of TGRF.
@@
First, the TGRF results in a higher performance. As shown in Tables I and II, the TGRF outperformed significantly in terms of the SR. Notably, Table I shows that the TGRF achieves an average SR of 94.091%, whereas the reward function without TGRF attains only 77.1% in the GST + HH Attn policy, marking a notable 17% increase.
@@
Hence, models employing the TGRF generally exhibit equal or superior performances in terms of the SR and ITR. This implies that TGRF demonstrates relatively high performance across various models. This signifies that TGRF effectively incorporates prior knowledge based on the role of the reward function, indicating resilience in the freezing robot problem [43]. As a result, it is evident that the robot demonstrates high SR by taking appropriate actions according to the situation.
@@
Second, even when combined with different reward functions without TGRF, TGRF exhibits high performance. In Tables I and II, for models utilizing trajectory prediction, $r_{pred}(s_{t})$ employed the previous paper’s reward function without TGRF. Nonetheless, it showcased superior performance compared to [13], proving its adaptability to other reward functions.
@@
Third, this leads to an enhanced recognition of human intent and collision avoidance. In Fig. 7a, the behavior of the robot when a reward function without TGRF was applied. The robot evidently struggles when confronted with a gathering crowd. Notably, in the test cases, the robot ventured into a crowd, resulting in unintended collisions with humans while attempting to navigate the crowd. Similar situations occurred in the other test cases. This means that the robot select aggressive or impolite behaviors, such as sidestepping to avoid human and unintentional collisions, or making risky decisions to reach a destination faster, resulting in collisions. This behavior reflects a deficiency in understanding the broader intentions of humans and imbalance of the reward functions.
@@
However, in Fig. 7b, the robot proactively positions itself behind the crowd before converging at a single point. This means that the values of $r_{pred}(s_{t})$, $r_{pot}(s_{t})$, and $r_{disc}(s_{t})$ related to human interactions were well balanced, enabling the robot to navigate effectively without colliding with individuals.
@@
Figure 8: Comparison of performance among four types of reward functions. The blue represents the reward function incorporating TGRF, while the orange corresponds to the reward function in [13], the green reflects [16], and the red signifies [31]. (a) denotes SR, (b) represents NT, (c) stands for PL, and (d) signifies ITR.
@@
This signifies that TGRF effectively incorporates prior knowledge, and its priorities are well integrated into the policy. It suggests that the algorithm’s performance can be further enhanced when TGRF is applied. Further evidence of this enhancement is reflected in the results in Tables I and II, where the average SR and standard deviation show similar or superior performances compared with previous iterations. In other test cases, we could observe that the robot selected a more secure and effective route than a faster and more dangerous route.

### v2

@@ 3.2.1 Formula and number of hyperparameters
TGRF leverages the characteristics of a normal distribution [47]. A normal distribution offers the advantage of being able to transform into various shapes using only two hyperparameters. This flexibility enhances the adaptability of the model, allowing adjustments to fit diverse prior knowledge and applying them to reward functions. Furthermore, it alleviates the burden on researchers in tuning hyperparameters, aiding in the swift identification of appropriate hyperparameters within a short timeframe.
@@
Figure 4: TGRF. The X-axis denotes the X-position in meters, Y-axis represents the Y-position in meters, and Z-axis indicates TGRF value when $w_{TGRF}=1$,$x_{TGRF}=dist(X,Y)$,$\mu_{TGRF}=0$,$\sigma_{TGRF}=2$.
@@
TGRF leverages the characteristics of a normal distribution. In (2), TGRF involves three hyperparameters: $w_{TGRF}$, representing the weight of TGRF; $\mu_{TGRF}$, denoting the mean; and $\rho_{TGRF}$, indicating the variance.
It incorporates a single variable $x_{TGRF}$. $C_{norm}$ ensures that the TGRF attains a maximum value of 1, irrespective of $\rho_{TGRF}$. This allows the scaling of the TGRF solely by $w_{TGRF}$, enabling researchers to intuitively control its maximum value, preventing the freezing robot problem [43], and balancing it with other reward values to attain the desired algorithmic performance.
$\rho_{TGRF}$ determines the transformability of TGRF. As $\lim_{\sigma_{TGRF}\to\infty}\mathcal{N}(x_{TGRF};\mu_{TGRF},\sigma_{TGRF})$, it takes on a constant form insensitive to changes in $x_{TGRF}$, while as $\lim_{\sigma_{TGRF}\to 0}\mathcal{N}(x_{TGRF};\mu_{TGRF},\sigma_{TGRF})$, it resembles an impulse function. This versatility enables the creation of diverse forms of TGRF, such as constant, linear, non-linear, Gaussian, etc., adaptable to specific objectives.
$\mu_{TGRF}$ determines the position of TGRF’s maximum value. Thus, allowing adjustments of the maximum value position according to the objectives requiring specific values to be reached. The variable $x_{TGRF}$ signifies the variable in N($x_{TGRF}$; $\mu_{TGRF}$,$\rho_{TGRF}$). As a result, it ultimately represents a shape similar to Fig. 4.
@@ 3.2.2 Transformability
Figure 5: Transformability of TGRF. The X-axis denotes the X-position in meters, Y-axis represents the Y-position in meters, and Z-axis indicates TGRF value. In (a), $w_{TGRF}=1$,$x_{TGRF}=dist(X,Y)$,$\mu_{TGRF}=0$,$\sigma_{TGRF}=5$. In (b), $w_{TGRF}=1$,$x_{TGRF}=dist(X,Y)$,$\mu_{TGRF}=0$,$\sigma_{TGRF}=5000$.
The reward functions described in Section II-B have different formulas and limited flexibility. These constraints lead to complexity in the reward functions and difficulty in determining optimal performance. Therefore, researchers must redesign and tune them to achieve better performance. However, the TGRF offers versatility in generating various shapes.
Fig. 5 illustrates the creation of different shapes using the same TGRF by simply adjusting $\rho_{TGRF}$. Fig. 5a shows a TGRF that generates a continuous Gaussian distribution, making it suitable for moving objects or humans by applying varying penalties. Fig. 5b shows the discrete column-like shape. This configuration is suitable for stationary reward functions and objects. This approach allows various shapes via adjusting only a hyperparameter ($\rho_{TGRF}$) without requiring an additional function. This significantly reduces researchers’ time and effort while enabling fine-tuning to match the specific characteristics of objects.
@@ 3.2.3 Reward function
Reward $r(s_{t},a_{t})$ is categorized into five types. First, $r_{goal}(s_{t})$ = 10 represents the reward when the robot successfully reaches its destination. Second, $r_{col}(s_{t})$ = -10 serves as a penalty incurred upon colliding with another individual. Third, $r_{disc}(s_{t})$ represents the penalty for entering a danger zone. Fourth, $r_{pot}(s_{t})$ corresponds to the reward/penalty contingent on the change in distance to the destination $S_{goal}$. Finally, $r_{pred}(s_{t})$ denotes the penalty invoked when entering a prediction.
Figure 6: TGRF applied to reward functions. The X-axis denotes the X-position in meters, Y-axis represents the Y-position in meters, and Z-axis indi-cates negative reward value. Central cylinder represents $r_{col}$, and surrounding distribution represents $r_{disc}(s_{t})$. Beyond $d_{disc}$, $r_{disc}(s_{t})$ becomes 0 ($s_{t}\notin S_{\text{danger zone}}$). In (a), $w_{TGRF}=8$,$x_{TGRF}=dist(X,Y)$,$\mu_{TGRF}=0$,$\sigma_{TGRF}=3$. In (b), $w_{TGRF}=10$,$x_{TGRF}=dist(X,Y)$, $\mu_{TGRF}=0$, $\sigma_{TGRF}=1000$.
The penalty $r_{disc}(s_{t})$ is imposed when the robot enters the danger zone ($d_{min}$ is within $d_{disc}$) determined by the nearest human distance, denoted as $d_{min}$. The formula used is as follows:
@@
In (3) and Fig. 6a, $r_{disc}(s_{t})$ is designed to prevent collisions with humans, following a Gaussian distribution with $\mu_{disc}$ = 0, whereby smaller values of $d_{min}$ (closer proximity to humans) result in a higher penalty. This design aligns with prior knowledge and exhibits a plausible and realistic shape.
$r_{disc}(s_{t})$ can be adjusted using only two hyperparameters. By tuning $\rho_{disc}$, researchers can regulate the breadth of the Gaussian penalty concerning the distance between humans and the robot. This enables the robot to react more sensitively or less sensitively to the distance from humans. $w_{disc}$ directly scales the reward function, establishing a linear correlation with $r_{disc}(s_{t})$, thereby enabling adjustment of the overall reward balance to prioritize driving tasks.
The potential reward $r_{pot}(s_{t})$ represents the reward associated with the potential field and is defined as follows:
@@
$r_{pot}(s_{t})$ plays a crucial role in guiding a robot toward its destination and mitigating the freezing robot problem [21]. However, high values of $r_{pot}(s_{t})$ can lead to increased collision rates in humans. Therefore, in Equation (4), by setting $x_{pot}=\mu_{pot}$, we aimed to maintain a constant TGRF regardless of $\mu_{pot}$ and $\rho_{pot}$. In addition, by applying $w_{pot}$ = 1.5, we ensured equal weighting for all $\Delta d$, aiming for uniformity across them.
@@
$r_{pred}(s_{t})$ is used only in models that employ trajectory predictions. $r_{pred}(s_{t})$ denotes the penalty value when the robot is positioned along the trajectory of the $i-th$ person. $\mathbbm{1}_{i}^{t+k}$ indicates whether the robot is in the predicted position of the $i-th$ person at time t+k or not. Thus, $r_{pred}(s_{t})$ takes the smallest penalty among all individual trajectory penalties that the robot takes. In our experiments, to demon-strate the performance enhancement even with different reward functions, we adopted $r_{pred}(s_{t})$ used in the GST + HH Attn from [13]. The final definition of the reward function is as follows:
@@
In summary, TGRF offers the distinct advantage of intuitively and efficiently modifying the reward function with fewer hyperparameters. This enables the robot to make rational decisions and reduces time-consuming fine-tuning tasks.
@@ 4.1 Experimental environment
Figure 7: Comparing scenarios with and without TGRF. Yellow circles represent robots, blue circles represent humans within the sensor range, red circles represent humans outside the sensor range, and orange circles in front of the blue circles indicate trajectories predicted by the GST.
We employed a 2D environmental simulator as in previous studies [13]. This simulator features a 12 × 12m space, with a 360° field of view and 5 m sensor range for LiDAR. A fixed number of humans (20) was used to represent a crowded setting.
@@
Both humans and robots were operated using holonomic kinematics to determine their velocities $(a_{t}=[v_{x},v_{y}]m/s)$ along the x- and y-axes. Holonomic kinematics refers to a state in which degrees of freedom can move independently without any constraints. This implies that robots and machines can move without limitations on their position or orientation. As a result, the action space of a robot is continuous, allowing both robots and humans to immediately achieve their desired speed within a time frame of $\Delta t$, assuming they operate within the maximum speed limit. Therefore, the positions of the humans and robots are continuously updated according to (7).
The robot has attributes such as a size of $\rho=0.3m$ and a maximum speed of $v_{max}$ = $1.0m/s$. Humans also have characteristics such as a size ranging from 0.3 to 0.5m and a maximum speed varying between 0.5 and 1.5m/s. In addition, the locations and destinations of the robot were randomized, and the destinations were not set excessively close. If the robot collides with an individual, reaches its destination, or exceeds the maximum time T, the episode is terminated, leading to the beginning of a subsequent episode.
Random seeding was applied during training, resulting in varying outcomes for each training episode. To handle the varying outcomes, multiple training runs were conducted with a total time step of 2×$10^{7}$ for DS-RNN and 1×$10^{7}$ for the other algorithms. The learning rate was set to 4×$10^{-5}$ for all policies. Subsequently, test data were acquired from 500 test episodes. The evaluation metrics applied to the test data included the success rate (SR), average navigation time (NT) in seconds, path length (PL) in meters for successful episodes, and intrusion time ratio (ITR).
@@ 4.2.1 Results in Different Navigation Methods and Environment
We compared the performance with the rewards (from Section III-B-3) when the TGRF was applied to the performance without its application (reward function from [13]). The experiment involved the application of five different navigation methods to the robots, as described in Section III-A-1. Table I shows the performance when individuals adhere to ORCA, whereas Table II outlines the performance when adhering to SF. The hyperparameters are set to $w_{disc}=0.25$, $x_{disc}=d_{min}$, $\mu_{disc}=0$,$\rho_{disc}=0.2$, $d_{disc}=0.5$, $w_{pot}=1.5$,$x_{pot}=\mu_{pot}$, $\mu_{pot}=0$,$\rho_{pot}=1000$. From Tables I, and II and Fig. 7, we can identify three impacts of the TGRF.
@@
First, the TGRF results in higher performance. As shown in Tables I and II, the TGRF significantly outperformed it in terms of SR. Notably, Table I shows that TGRF achieves an average SR of 94.091%, whereas the reward function without TGRF attains only 77.1% under the GST + HH Attn policy, marking a notable 17% increase.
@@
Hence, models employing the TGRF generally exhibit equal or superior performance in terms of SR and ITR. This implies that the TGRF demonstrates relatively high performance across various models. This signifies that the TGRF effectively incorporates prior knowledge based on the role of the reward function, indicating resilience in the freezing robot problem [43]. Consequently, it is evident that the robot demonstrates high SR by taking appropriate actions according to the situation.
@@
Second, even when combined with different reward functions without the TGRF, the TGRF exhibits high performance. In Tables I and II, for models utilizing trajectory prediction, $r_{pred}(s_{t})$ employs the reward function of a previous study without the TGRF. Nonetheless, it showed superior performance compared with [13], proving its adaptability to other reward functions.
@@
Third, this leads to enhanced recognition of human intent and collision avoidance. Fig. 7a shows the behavior of the robot when a reward function without the TGRF was applied. The robot struggles when confronted with crowds. Notably, in the test cases, the robot ventured into the crowd, resulting in unintended collisions with humans while attempting to navigate the crowd. Similar situations were observed in the other test cases. This means that the robot selects aggressive or impolite behaviors, such as sidestepping, to avoid human and unintentional collisions, or making risky decisions to reach a destination faster, resulting in collisions. This behavior reflects a deficiency in understanding the broader intentions of humans and an imbalance in reward functions.
@@
However, in Fig. 7b, the robot proactively positioned itself behind the crowd before converging at a single point. This means that the values of $r_{pred}(s_{t})$, $r_{pot}(s_{t})$, and $r_{disc}(s_{t})$ related to human interactions were well balanced, enabling the robot to navigate effectively without colliding with individuals.
@@
Figure 8: Comparison of performance among four types of reward functions. Blue line represents the reward function incorporating TGRF, while or-ange line corresponds to the reward function in [13], green one reflects [16], and red signifies [31]. (a) denotes SR, (b) represents NT, (c) stands for PL, and (d) signifies ITR.
@@
This signifies that the TGRF effectively incorporates prior knowledge and that its priorities are well integrated into the policy. This suggests that the performance of the algorithm can be further enhanced when the TGRF is applied. Further evidence of this enhancement is reflected in the results in Tables I and II, where the average SR and standard deviation show similar or superior performances compared with previous iterations. In the other test cases, we observed that the robot selected a more secure and effective route than a faster and more dangerous route.